In solving the following system of equations, what is the value of the discriminant in the resulting equation? y = x2 - 2x + 1 and y = 1/4x^2 + 1/2
Answer by Alan3354(69443)

what is the value of the discriminant in the resulting equation?
y = x^2 - 2x + 1
Disc = b^2 - 4ac
a = 1, b = -2, c = 1
Disc = (-2)^2 - 4*1*1 = 4-4
Disc = 0
----------------
and y = 1/4x^2 + 1/2
a = 1/4, b = 0, c = 1/2